he width of a rectangle is 2 less than one half its length. if perimeter is 86 cm whats the equation
Answer by Cromlix(4381)

Hi there,
Make length = x
width = x/2 - 2
Perimeter = 2(length) + 2(width)
86 = 2(x) + 2(x/2 - 2)
86 = 2x + x - 4
Collect like terms:
2x + x = 86 + 4
3x = 90
x = 30
Length = 30cm
Width = 13cm
Hope this helps :-)